# DynFaceRestore: Balancing Fidelity and Quality in Diffusion-Guided Blind Face Restoration with Dynamic Blur-Level Mapping and Guidance

### v1

@@ Abstract
Blind Face Restoration aims to recover high-fidelity, detail-rich facial images from unknown degraded inputs, presenting significant challenges in preserving both identity and detail. Pre-trained diffusion models have been increasingly used as image priors to generate fine details. Still, existing methods often use fixed diffusion sampling timesteps and a global guidance scale, assuming uniform degradation. This limitation and potentially imperfect degradation kernel estimation frequently lead to under- or over-diffusion, resulting in an imbalance between fidelity and quality. We propose DynFaceRestore, a novel blind face restoration approach that learns to map any blindly degraded input to Gaussian blurry images. By leveraging these blurry images and their respective Gaussian kernels, we dynamically select the starting timesteps for each blurry image and apply closed-form guidance during the diffusion sampling process to maintain fidelity. Additionally, we introduce a dynamic guidance scaling adjuster that modulates the guidance strength across local regions, enhancing detail generation in complex areas while preserving structural fidelity in contours. This strategy effectively balances the trade-off between fidelity and quality. DynFaceRestore achieves state-of-the-art performance in both quantitative and qualitative evaluations, demonstrating robustness and effectiveness in blind face restoration.
@@ 1 Introduction
Figure 2: Using DifFace [41] as an example, let $RM$ represent DifFace’s restoration model and $p_{\Theta}$ the diffusion model. In the t-SNE plot, green and red points denote the features of $x_{400}^{0}$ sampled from $p_{\Theta}(x_{400}^{0}|\acute{y})$ and $p_{\Theta}(x_{400}^{0}|x_{0})$, respectively. Here, $\acute{y}$, the LQ image restored by $RM$, is used for diffusion guidance. DifFace initiates the diffusion process at a fixed $t=400$, resulting in under-diffusion (right) for severely degraded LQ images and just enough diffusion for mildly degraded LQ images (left). This underscores the importance of selecting an appropriate starting step.
@@ 3.2 DynFaceRestore Sampling
Rather than directly tackling the challenging BFR task, our DynFaceRestore method reframes it as a Gaussian deblurring problem by transforming the blindly degraded input image ${y}$ into a Gaussian-blurred version $\acute{y}$ and applying a modified DPS-based guidance for restoration. This allows us to represent $\mathcal{A}(x_{t}^{0})$ as $k_{t}\otimes x_{t}^{0}$, where $k$ is a Gaussian blur kernel with standard deviation $std$. Consequently, our guided diffusion sampling process can be reformulated as follows:
@@
which are derived from diffusion model. However, perfect kernel estimation is often impractical; to reduce kernel mismatch, we also refine the $\mathcal{A}$ during diffusion sampling. Since our $\mathcal{A}$ is strictly a Gaussian blur kernel, we only need to update the estimated $std$ by using the following equation.
@@
5:$t_{start}=DSST(\hat{std^{*}})$;
6:$x_{t_{start}}=\sqrt{\bar{\alpha}_{t_{start}}}\acute{y}+\sqrt{1-\bar{\alpha}_{t_{start}}}\epsilon,\epsilon\sim N(0,1)$;
7:$std_{t_{start}}=\hat{std^{*}}$;
@@ 4.1.1 Estimation of ${std^{*}}$
During inference, the ground truth $x$ is unavailable, necessitating an alternative approach for estimating ${std^{*}}$ instead of using Eq. 19. To achieve this, we design a learnable network (SE) to generate an estimate $\hat{{std^{*}}}$ from the input $y$. SE first produces an intermediate output $\hat{y}^{\prime}$ that approximates $\tilde{x}$, and then, based on $\hat{y}^{\prime}$ derives $\hat{{std^{*}}}$. In summary, $[{\hat{std^{*}},\hat{y}^{\prime}}]=SE(y)$. The training objective function for SE is defined as follows:
$\begin{split}L_{SE}=\mathbb{D}(\hat{std^{*}},std^{*})+\gamma_{std}\mathbb{D}(\hat{y}^{\prime},\tilde{x})),\end{split}$ (10)
where $\mathbb{D}$ is the L1 distance and $\gamma_{std}$ is a weighting factor for balancing. Additionally, ${std^{*}}$ and ${\tilde{x}}$ are the training labels derived from Sec. 4.1. Note that $\hat{y}^{\prime}$, predicted by the SE, maps the unknown degraded input image ${y}$ to an intermediate blurred version, serving as an alternative to the Gaussian-blurred counterpart $\acute{y}$ from Eq. 8. However, $\hat{y}^{\prime}$ does not perfectly conform to the assumed Gaussian form and is unsuitable as Gaussian-blurred guidance in the subsequent diffusion sampling process. Thus, in our final implementation, we use $\acute{y}$ instead of $\hat{y}^{\prime}$, estimating $\acute{y}$ via Eq. 8, with $\hat{{std^{*}}}$ replaced by $\hat{{std^{*}}}$. Due to space constraints, a detailed analysis of the differences between $\acute{y}$ and $\hat{y}^{\prime}$, along with the design details of the SE, is provided in the Supplementary.
@@ 4.2 Dynamic Starting Step Lookup Table
Based on the estimated blur level, $\hat{{std^{*}}}$, we develop a strategy to determine the starting timesteps for inserting guidance $\acute{y}$ into DM to prevent under- and over-diffusion. The idea is illustrated in Fig. 4. Assume $x_{0}\equiv x$ represents an HQ image and $\tilde{y}_{0}^{std}=k^{std}\otimes x$ is its blurred counterpart. When both images undergo the same forward degradation process defined in a standard diffusion model, they will statistically converge after a specific timestep $t$, meaning $x_{t}\cong\tilde{y}_{t}^{std}$. The timestep $t$ then serves as a suitable insertion point for a blurry observation $\tilde{y}_{0}^{std}$ with blur level $std$. To identify the optimal starting timestep $t_{std}$ for each standard deviation, we use the following equation:
[…]
where $tol$ represents the maximum tolerance, $\textbf{X}_{t}$ and $\tilde{\textbf{Y}}_{t}^{std}$ denote the expected values of $x_{t}$ and $\tilde{y}_{t}^{std}$ in a training set.
@@ 4.4 Multiple Mapping and Guidance
Up to this point, we have employed only a single fidelity guidance, $\acute{y}$, and relied on the diffusion model (DM) to recover the missing high-frequency components (i.e., the difference between $x$ and $\acute{y}$) for restoration. However, while the DM effectively enhances image quality, it often sacrifices fidelity; conversely, the restoration model (RM) maintains fidelity but yields lower quality enhancement. To balance the strengths of DM and RM, we propose an optional add-on module that leverages multiple guidance sources. Specifically, once we obtain the estimated $\hat{{std^{*}}}$ from SE, we generate three Gaussian-blurred guidance images $\acute{y}^{i\in[1,2,3]}$ by varying $std$ values. Formally,
@@
Since $\acute{y}^{1}$ retains more trustable low-frequency structural information, while $\acute{y}^{2}$ and $\acute{y}^{3}$ provide more high-frequency details but with lower confidence, we prioritize adjustments derived from $\acute{y}^{1}$ by assigning it a higher weight, while assigning smaller weights to the guidance from $\acute{y}^{2}$ and $\acute{y}^{3}$. Additionally, the standard deviation of $k^{i\in[1,2,3]}$ are continuously adjusted throughout the diffusion process, as described in Eq. 7, which mitigates the impact of kernel mismatch and thereby enhances the fidelity of the final result. Nonetheless, one remaining issue is the selection of hyperparameters, including the number of guidance and the settings of $std$ values, which can be explored in future work.
@@ 5.2 Comparisons to SOTA methods
We conduct comprehensive quantitative and qualitative comparisons of our proposed DynFaceRestore with the following state-of-the-art BFR methods: $GAN-based$ (GPEN [40], GFPGAN [22]), $Codebook-based$ (RestoreFormer [36], CodeFormer [43], DAEFR [33]) and $DiffusionModel(DM)-based$ (DiffBIR [26], PGDiff [38], DifFace [41], DR2 [37] and 3Difffusion [27]).
@@ A.1 Dynamic Blur-Level Mapping
Here, $\xi$ represents the error tolerance, $x$ denotes the high-quality (HQ) ground truth, and $\otimes$ signifies the convolution operation. The search space for the standard deviation range, $[std_{min},std_{max}]$, is defined between 0.1 and 15.0 with an interval of 0.1. To solve the optimization problem in Eq. 19, we employ a brute-force method.
@@
Figure 8: The SE consists of $TM$ and a $SDE$, predicting the Gaussian blur level corresponding to the degraded input based on intermediate Gaussian blur image $\hat{y}^{\prime}$.
Table 4: Architecture of $SDE$.
@@
Conv2d: kernel size: 3$\times$3; stride: 1 3 $\rightarrow$ 64
@@
Conv2d: kernel size: 3$\times$3; stride: 1 64 $\rightarrow$ 64
@@
Conv2d: kernel size: 3$\times$3; stride: 2 64 $\rightarrow$ 128
@@
Conv2d: kernel size: 3$\times$3; stride: 1 128 $\rightarrow$ 128
@@
Conv2d: kernel size: 3$\times$3; stride: 2 128 $\rightarrow$ 256
@@
Conv2d: kernel size: 3$\times$3; stride: 1 256 $\rightarrow$ 256
@@
$\begin{split}L_{SE}=\mathbb{D}(\hat{std^{*}},std^{*})+\gamma_{std}\mathbb{D}(\hat{y}^{\prime},\tilde{x}),\end{split}$ (22)
@@ A.2 Dynamic Starting Step Look-up Table
[…]
where $tol$ represents the maximum tolerance, $\textbf{X}_{t}$ and $\tilde{\textbf{Y}}_{t}^{std}$ denote the expected values of $x_{t}$ and $\tilde{y}_{t}^{std}$ in a training set. We set $tol$ to $1\times 10^{-3}$ to construct the Dynamic Starting Step Look-up Table (DSST), which pairs each standard deviation $std$ with its corresponding starting step $t_{std}$. Given the Gaussian-blur image $\acute{y}$, the estimated standard deviation $\hat{std^{*}}$ from SE is equalized and serves as a key to retrieve the corresponding starting step from the DSST.
@@
Conv2d: kernel size: 3$\times$3 6 $\rightarrow$ 64
@@
Conv2d: kernel size: 3$\times$3 64 $\rightarrow$ 64
@@
Conv2d: kernel size: 3$\times$3 64 $\rightarrow$ 3
@@ A.3 Dynamic Guidance Scale Adjuster
15:$\quad\quad\quad\quad\quad\quad+\quad DISTS(x_{t-1}^{0},x_{0})$;
@@
where $\gamma_{i}$ are the weighted factors of the four subbands (LL, LH, HL, HH) decomposed by Stationary Wavelet Transformation ($SWT$) [15, 21], $x_{0}$ is the learning target, and $x_{t-1}^{0}$ is the HQ prediction based on $x_{t-1}$ at timestep $t-1$. Here, $\mathbb{D}$ and $DISTS$ [8, 21] are the L1 reconstruction loss and the perceptual loss, respectively.
@@ A.4 Inference
6:$\acute{y}^{i\in[1,2,3]}=[\acute{y},k^{2}\otimes RM(y),k^{3}\otimes RM(y)]$
7:$t_{start}^{i\in[1,2,3]}=DSST(\hat{std^{*}}^{i\in[1,2,3]})$;
8:$std_{t_{start}}^{i\in[1,2,3]}=\hat{std^{*}}^{i\in[1,2,3]}$;
9:$x_{t_{start}}=\sqrt{\bar{\alpha}_{t_{start}^{1}}}\acute{y}^{1}+\sqrt{1-\bar{\alpha}_{t_{start}^{1}}}\epsilon,\epsilon\sim N(0,1)$;
@@ B.1 Kernel Mismatch
Figure 10: The frequency responses of actual Gaussian blurry images, $\acute{y}$, and $\hat{y}^{\prime}$. Here, $\hat{y}^{\prime}$, the output of $TM$ (refer to Fig. 8), exhibits significant differences in the high-frequency components compared to actual Gaussian-blurred images. In contrast, $\acute{y}$, estimated using Eq. 17, closely matches the frequency response of the actual Gaussian-blurred image.
@@ B.3 Multiple Guidance
Table 7: Ablations of the different numbers of guidance in CelebA-Test dataset. ##\## denotes the numbers of guidance and $\lambda^{i}$ are the weights of each guidance. The best performance is highlighted with bold.
@@ B.4 Different Starting Steps
By setting up the output of DBLM as a Gaussian-blurred image based on the input degradation severity, we can leverage this property to identify the optimal timestep ($t^{*}$), as outlined in Sec. 4.2 using “Dynamic Starting Step Lookup Table”. Note that $t^{*}$ is automatically determined for different datasets. To demonstrate the effectiveness of our approach, we conduct experiments with different starting timesteps, as shown in Tab. 8. Larger timesteps ($>t^{*}$) result in a loss of information from the guidance observation, leading to reduced fidelity. Conversely, smaller timesteps ($<t^{*}$) provide insufficient iterations to recover fine details, thereby diminishing the quality of the reconstruction.

### v2

@@ Abstract
Blind Face Restoration aims to recover high-fidelity, detail-rich facial images from unknown degraded inputs, presenting significant challenges in preserving both identity and detail. Pre-trained diffusion models have been increasingly used as image priors to generate fine details. Still, existing methods often use fixed diffusion sampling timesteps and a global guidance scale, assuming uniform degradation. This limitation and potentially imperfect degradation kernel estimation frequently lead to under- or over-diffusion, resulting in an imbalance between fidelity and quality. We propose DynFaceRestore, a novel blind face restoration approach that learns to map any blindly degraded input to Gaussian blurry images. By leveraging these blurry images and their respective Gaussian kernels, we dynamically select the starting timesteps for each blurry image and apply closed-form guidance during the diffusion sampling process to maintain fidelity. Additionally, we introduce a dynamic guidance scaling adjuster that modulates the guidance strength across local regions, enhancing detail generation in complex areas while preserving structural fidelity in contours. This strategy effectively balances the trade-off between fidelity and quality. DynFaceRestore achieves state-of-the-art performance in both quantitative and qualitative evaluations, demonstrating robustness and effectiveness in blind face restoration. Project page at Link
@@ 1 Introduction
Figure 2: Using DifFace [41] as an example, let RMRM represent DifFace’s restoration model and $p_{\Theta}$ the diffusion model. In the t-SNE plot, green and red points denote the features of $x_{400}^{0}$ sampled from $p_{\Theta}(x_{400}^{0}|\acute{y})$ and $p_{\Theta}(x_{400}^{0}|x_{0})$, respectively. Here, $\acute{y}$, the LQ image restored by $RM$, is used for diffusion guidance. DifFace initiates the diffusion process at a fixed $t=400$, resulting in under-diffusion (right) for severely degraded LQ images and just enough diffusion for mildly degraded LQ images (left). This underscores the importance of selecting an appropriate starting step.
@@ 3.2 DynFaceRestore Sampling
Rather than directly tackling the challenging BFR task, our DynFaceRestore method reframes it as a Gaussian deblurring problem by transforming the blindly degraded input image y{y} into a Gaussian-blurred version $\acute{y}$ and applying a modified DPS-based guidance for restoration. This allows us to represent $\mathcal{A}(x_{t}^{0})$ as $k_{t}\otimes x_{t}^{0}$, where $k$ is a Gaussian blur kernel with standard deviation stdstd. Consequently, our guided diffusion sampling process can be reformulated as follows:
@@
which are derived from diffusion model. However, perfect kernel estimation is often impractical; to reduce kernel mismatch, we also refine the $\mathcal{A}$ during diffusion sampling. Since our $\mathcal{A}$ is strictly a Gaussian blur kernel, we only need to update the estimated stdstd by using the following equation.
@@
[…]
@@ 4.1.1 Estimation of $std^{*}$}
During inference, the ground truth $x$ is unavailable, necessitating an alternative approach for estimating $std^{*}$} instead of using Eq. 19. To achieve this, we design a learnable network (SE) to generate an estimate $\hat{{std^{*}}}$ from the input $y$. SE first produces an intermediate output $\hat{y}^{\prime}$ that approximates $\tilde{x}$, and then, based on $\hat{y}^{\prime}$ derives $\hat{{std^{*}}}$. In summary, ${\hat{std^{*}},\hat{y}^{\prime}}]=SE(y)$. The training objective function for SE is defined as follows:
[…]
where $\mathbb{D}$ is the L1 distance and $\gamma_{std}$ is a weighting factor for balancing. Additionally, $std^{*}$} and $\tilde{x}$} are the training labels derived from Sec. 4.1. Note that $\hat{y}^{\prime}$, predicted by the SE, maps the unknown degraded input image y{y} to an intermediate blurred version, serving as an alternative to the Gaussian-blurred counterpart $\acute{y}$ from Eq. 8. However, $\hat{y}^{\prime}$ does not perfectly conform to the assumed Gaussian form and is unsuitable as Gaussian-blurred guidance in the subsequent diffusion sampling process. Thus, in our final implementation, we use $\acute{y}$ instead of $\hat{y}^{\prime}$, estimating $\acute{y}$ via Eq. 8, with $\hat{{std^{*}}}$ replaced by $\hat{{std^{*}}}$. Due to space constraints, a detailed analysis of the differences between $\acute{y}$ and $\hat{y}^{\prime}$, along with the design details of the SE, is provided in the Supplementary.
@@ 4.2 Dynamic Starting Step Lookup Table
Based on the estimated blur level, $\hat{{std^{*}}}$, we develop a strategy to determine the starting timesteps for inserting guidance $\acute{y}$ into DM to prevent under- and over-diffusion. The idea is illustrated in Fig. 4. Assume $x_{0}\equiv x$ represents an HQ image and $\tilde{y}_{0}^{std}=k^{std}\otimes x$ is its blurred counterpart. When both images undergo the same forward degradation process defined in a standard diffusion model, they will statistically converge after a specific timestep $t$, meaning $x_{t}\cong\tilde{y}_{t}^{std}$. The timestep $t$ then serves as a suitable insertion point for a blurry observation $\tilde{y}_{0}^{std}$ with blur level stdstd. To identify the optimal starting timestep $t_{std}$ for each standard deviation, we use the following equation:
$t_{std}=\underset{t}{argmin}\quad(log(\textbf{X}_{t})-log(\tilde{\textbf{Y}}_{t}^{std})\leq tol),$ (11)
where toltol represents the maximum tolerance, $\textbf{X}_{t}$ and $\tilde{\textbf{Y}}_{t}^{std}$ denote the expected values of $x_{t}$ and $\tilde{y}_{t}^{std}$ in a training set.
@@ 4.4 Multiple Mapping and Guidance
Up to this point, we have employed only a single fidelity guidance, $\acute{y}$, and relied on the diffusion model (DM) to recover the missing high-frequency components (i.e., the difference between $x$ and $\acute{y}$) for restoration. However, while the DM effectively enhances image quality, it often sacrifices fidelity; conversely, the restoration model (RM) maintains fidelity but yields lower quality enhancement. To balance the strengths of DM and RM, we propose an optional add-on module that leverages multiple guidance sources. Specifically, once we obtain the estimated $\hat{{std^{*}}}$ from SE, we generate three Gaussian-blurred guidance images $\acute{y}^{i\in[1,2,3]}$ by varying stdstd values. Formally,
@@
Since $\acute{y}^{1}$ retains more trustable low-frequency structural information, while $\acute{y}^{2}$ and $\acute{y}^{3}$ provide more high-frequency details but with lower confidence, we prioritize adjustments derived from $\acute{y}^{1}$ by assigning it a higher weight, while assigning smaller weights to the guidance from $\acute{y}^{2}$ and $\acute{y}^{3}$. Additionally, the standard deviation of $k^{i\in[1,2,3]}$ are continuously adjusted throughout the diffusion process, as described in Eq. 7, which mitigates the impact of kernel mismatch and thereby enhances the fidelity of the final result. Nonetheless, one remaining issue is the selection of hyperparameters, including the number of guidance and the settings of stdstd values, which can be explored in future work.
@@ 5.2 Comparisons to SOTA methods
We conduct comprehensive quantitative and qualitative comparisons of our proposed DynFaceRestore with the following state-of-the-art BFR methods: GANbasedGAN-based (GPEN [40], GFPGAN [22]), CodebookbasedCodebook-based (RestoreFormer [36], CodeFormer [43], DAEFR [33]) and DiffusionModel(DM)basedDiffusionModel(DM)-based (DiffBIR [26], PGDiff [38], DifFace [41], DR2 [37] and 3Difffusion [27]).
@@ A.1 Dynamic Blur-Level Mapping
Here, $\xi$ represents the error tolerance, $x$ denotes the high-quality (HQ) ground truth, and $\otimes$ signifies the convolution operation. The search space for the standard deviation range, $std_{min},std_{max}]$, is defined between 0.1 and 15.0 with an interval of 0.1. To solve the optimization problem in Eq. 19, we employ a brute-force method.
@@
Figure 8: The SE consists of $TM$ and a SDESDE, predicting the Gaussian blur level corresponding to the degraded input based on intermediate Gaussian blur image $\hat{y}^{\prime}$.
Table 4: Architecture of SDESDE.
@@
Conv2d: kernel size: $\times3$; stride: 1 3 $\rightarrow$ 64
@@
Conv2d: kernel size: $\times3$; stride: 1 64 $\rightarrow$ 64
@@
Conv2d: kernel size: $\times3$; stride: 2 64 $\rightarrow$ 128
@@
Conv2d: kernel size: $\times3$; stride: 1 128 $\rightarrow$ 128
@@
Conv2d: kernel size: $\times3$; stride: 2 128 $\rightarrow$ 256
@@
Conv2d: kernel size: $\times3$; stride: 1 256 $\rightarrow$ 256
@@
[…]
@@ A.2 Dynamic Starting Step Look-up Table
$t_{std}=\underset{t}{argmin}\qquad(log(\textbf{X}_{t})-log(\tilde{\textbf{Y}}_{t}^{std})\leq tol),$ (23)
where toltol represents the maximum tolerance, $\textbf{X}_{t}$ and $\tilde{\textbf{Y}}_{t}^{std}$ denote the expected values of $x_{t}$ and $\tilde{y}_{t}^{std}$ in a training set. We set toltol to $1\times 10^{-3}$ to construct the Dynamic Starting Step Look-up Table (DSST), which pairs each standard deviation stdstd with its corresponding starting step $t_{std}$. Given the Gaussian-blur image $\acute{y}$, the estimated standard deviation $\hat{std^{*}}$ from SE is equalized and serves as a key to retrieve the corresponding starting step from the DSST.
@@
Conv2d: kernel size: $\times3$ 6 $\rightarrow$ 64
@@
Conv2d: kernel size: $\times3$ 64 $\rightarrow$ 64
@@
Conv2d: kernel size: $\times3$ 64 $\rightarrow$ 3
@@ A.3 Dynamic Guidance Scale Adjuster
[…]
@@
where $\gamma_{i}$ are the weighted factors of the four subbands (LL, LH, HL, HH) decomposed by Stationary Wavelet Transformation ($SWT$) [15, 21], $x_{0}$ is the learning target, and $x_{t-1}^{0}$ is the HQ prediction based on $x_{t-1}$ at timestep t1t-1. Here, $\mathbb{D}$ and $DISTS$ [8, 21] are the L1 reconstruction loss and the perceptual loss, respectively.
@@ A.4 Inference
[…]
@@ B.1 Kernel Mismatch
Figure 10: The frequency responses of actual Gaussian blurry images, $\acute{y}$, and $\hat{y}^{\prime}$. Here, $\hat{y}^{\prime}$, the output of TMTM (refer to Fig. 8), exhibits significant differences in the high-frequency components compared to actual Gaussian-blurred images. In contrast, $\acute{y}$, estimated using Eq. 17, closely matches the frequency response of the actual Gaussian-blurred image.
@@ B.3 Multiple Guidance
Table 7: Ablations of the different numbers of guidance in CelebA-Test dataset. #\# denotes the numbers of guidance and $\lambda^{i}$ are the weights of each guidance. The best performance is highlighted with bold.
@@ B.4 Different Starting Steps
By setting up the output of DBLM as a Gaussian-blurred image based on the input degradation severity, we can leverage this property to identify the optimal timestep ($t^{*}$), as outlined in Sec. 4.2 using “Dynamic Starting Step Lookup Table”. Note that $t^{*}$ is automatically determined for different datasets. To demonstrate the effectiveness of our approach, we conduct experiments with different starting timesteps, as shown in Tab. 8. Larger timesteps ($t^{*}$) result in a loss of information from the guidance observation, leading to reduced fidelity. Conversely, smaller timesteps ($t^{*}$) provide insufficient iterations to recover fine details, thereby diminishing the quality of the reconstruction.